A water tank with vertical sides has a

horizontal base in the shape of a rectangle with semi-circular ends. The total inside length of the tank is 7 m, its width 4m and its height 2 m. Calculate:

(a) the surface area of the vertical walls of the

tank in m2?

(b) the area of the base in m2

(c) the number of liters of water in the tank when the depth of water is 1.56m


1
Expert's answer
2022-05-25T15:08:40-0400

semi-circular ends:

"r = 4\/2=2" r = 4/2=2

rectangle

"width=7-2r=3"

(a) the surface area of the vertical walls of the tank in m2?

"(2\pi r+2*3)*2=(4 \pi+6)*2=8 \pi+12=37.13"

(b) the area of the base in m2

"3*4+\pi r^2 = 12+4\pi=24.57"

(c) the number of liters of water in the tank when the depth of water is 1.56m

"(12+4\pi)*1.56 = 38.323 m^3 = 38323 liters"